OpenStudy (anonymous):

Solve. sqrt6*[sin^2(x)-cos^2(x)] , sin2x=sqrt3/3 , -3PI/4

OpenStudy (dumbcow):

i don't get the equation, why is there a comma?

OpenStudy (anonymous):

Just to separate the conditions.

OpenStudy (anonymous):

\[\sqrt{6}*(\sin^2x-\cos^2x) . \]

OpenStudy (anonymous):

\[\sin2x=_{3}^{\sqrt{3}}\]

OpenStudy (dumbcow):

thats not an equation though however, sin^ -cos^2 = -cos(2x)

OpenStudy (anonymous):

That I know. But I need to get sin or cos from the sin2x and then plug them into the first equation to get the value of x.

OpenStudy (dumbcow):

ok gotcha, so you need to evaluate the expression given those conditions :)

OpenStudy (anonymous):

Yes. :)

OpenStudy (campbell_st):

draw a diagram and find the side marked a  \[x = \sqrt{6}\] \[\cos(2x) = \sqrt{6}/3\] so the problem is \[\sqrt{6}((\sqrt{3}/3)^2 - (\sqrt{6}/3)^2) = - 3\sqrt{6}/9 = \sqrt{6}/3\]